A long distance runner runs 4 miles south and then 6 miles east. How far is the runner from the starting point?
Answer by addingup(3677)

√(a^2+b^2) = √c^2
√(16+36) = √c^2
√52 = √c^2 take the square root of both sides:
7.21 = c
the runner is 7.21 miles from the starting point